ABC Ltd. is considering a system that will cost $750,000. The new system will be placed in CCA class 39 (CCA rate is 25%), has a useful life of six years, and an expected resale value of $100,000. The company has a tax rate of 40% and an opportunity cost of capital of 15%. What is the present value of the CCA tax shield?
Answer #1

Answer :-

Value of asset after 6 years :-

750000*(0.75)6 = 133484$

Note : the cost is multiplied With 0.75 because if cost is Re 1 then depriciation would be 0.25 so the post depriciation cost would be 0.75.

In this case there would be loss as the resale value would be $100000. Therefore the loss is : (133484-100000)$ = 33484$.

The loss would also provide tax shield to the company.

depriciation each year would be :-

Year cost price /post depriciation depriciation depriciation tax shield @40% present value @15%
0 750000
1 187500 75000 65217
2 562500 140625 56250 42533
3 421875 105469 42188 27739
4 316406 79102 31641 18091
5 237304 59326 23730 11798
6 177978 44494 17798 7695
Total 173073

Tax shield on loss on resale of asset = 33484*0.40*(1/1.156) =$5790

The total tax shield would be = 173073+5790 = $178863